OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN:what is the next term in the multiplication pattern that begins 13,39,117? A.1521 B.351 C.195 D.169

OpenStudy (anonymous):

@cwrw238 @campbell_st

OpenStudy (anonymous):

117*3=?

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

here first term a=39 common ratio r=39/13=3 it is a G.P \[tn=ar ^{n-1}\] \[t4=(13)3^{4-1}=13*3^3=13*3*3*3=?\]

OpenStudy (anonymous):

its 351 thank you so much you are very smart!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

what are the first four terms in the multiplication pattern in which the first term is 5 and then each term is multiplied by 6?

OpenStudy (anonymous):

5,30,180,1080

OpenStudy (anonymous):

YAYY THANKS

OpenStudy (anonymous):

here you multiply successively by 6

OpenStudy (anonymous):

DESCRIBE THE MULTIPLICATION PATTERN AND FIND THE NEXT TWO TERMS. 25; 100;400;1600;6400

OpenStudy (anonymous):

100/25=4

OpenStudy (anonymous):

here you multiply successively by 4

OpenStudy (anonymous):

LOL THANKS THAT WAS THE LAST ONE BYE :)

OpenStudy (anonymous):

you're welcome.